# Exploiting synergies between JWST and cosmic 21-cm observations to uncover star formation in the early Universe

### v1

@@ 6 Conclusions
    The star-formation efficiency model, as illustrated in Figure 1, consists of four parameters: $\tilde{f}_{\star,\mathrm{II}}$ scales the mangitude of the SFE, $M_{0}$ sets the halo mass below which the curve flattens, $\alpha_{\star}$ is the power law slope of the $M_{h}$-dependence, and $\beta_{\star}$ controls the redshift-dependence anchored to $z=6$. We find an interesting non-monotonic evolution of the SFE with redshift, with the SFE parameters almost entirely constrained by UVLF data (other data sets providing much weaker constraints). We split the UVLF data into two redshift bins $z=7\,$ 1010\,1010 and $z=11\,$ 14.514.5\,14.514.5 with $z=6$ data included in both bins to isolate the evolution parameter $\beta_{\star}$. We find a slow evolution of SFE at $z\lesssim 10$ with $\langle\beta_{\star}\rangle=1.4$ (and the posterior saturating to zero, consistent with no evolution), and then a rapid evolution at $z\gtrsim 11$ with $\langle\beta_{\star}\rangle=3.6$. The joint fit combining both redshift bins gives a maximum posterior estimate $\langle\beta_{\star}\rangle=3.0$. Thus, we find consistency with pre-JWST HST-calibrated results at low redshifts (Sun & Furlanetto, 2016; Tacchella et al., 2018), and post-JWST enhanged SFE models at high redshifts (in particular the feedback-free burst model of Dekel et al., 2023; Li et al., 2024). The slope of the SFE with halo mass $\langle\alpha_{\star}\rangle\approx 0.9$ is consistent across the two redshift bins at $M_{h}\gtrsim 10^{10}\,\mathrm{M}_{\odot}$, in agreement with the findings of Tacchella et al. (2018). Futhermore, under our composite model, we infer that DM halos of mass $M_{h}=10^{10}\,\mathrm{M}_{\odot}$ have a SFE of 222\,2 $\,3\%$ at $z\lesssim 10$, $12\%$ at $z=12$ and $26\%$ at $z=15$. Finally, we derive for the first time, functional posteriors for SFE in the low and high-redshift regime using the largest available dataset of galaxy observations.
@@
    The radio efficiency of early galaxies, defined relative to those in the local Universe in Equation 19, is constrained by a combination of CRB data and HERA data (with a weak contribution from SARAS 3). CRB data disfavours high radio efficiency of galaxies by their direct contribution to the diffuse background seen today. HERA further disfavours strong radio sources since they would boost the amplitude of 21-cm signal fluctuations beyond observed upper limits. The joint analysis results in an upper limit $f_{r}<11.7\,(170.3)$ at $68\%$ ($95\%$) confidence levels. We further rule out $f_{r}\gtrsim 1650$ at $5\sigma$ confidence level. At $1\sigma$, this is comparable to the constraints obtained by Sikder et al. (2024b) using radio source clustering data from VLA and ATCA (not included here).
@@
    We find that the analysis without UVLF data provides a weaker inference on $f_{X}$ and $f_{r}$, due to their degeneracy with the SFE of galaxies in the calculation of the respective diffuse backgrounds. For example, joint constraints without UVLFs disfavour $f_{X}\gtrsim 193$ and $f_{r}\gtrsim 2280$ at $2\sigma$. By including UVLF data and thus informing the model of SFE in observable galaxies (with small galaxies remaining poorly constrained), we break this degeneracy and tighten the constraints by a factor of $\sim 20$.
In conclusion, we have shown the critical role of UVLFs in our understanding of the Cosmic Dawn and Epoch of Reionization: they set an upper limit on the minimum mass of star-forming halos, constrain the star-formation efficiency of galaxies across cosmic epochs, and thus also break degeneracies with other astrophysical parameters constrained by different wavelength probes. In the future, we will extend this analysis to include other reionization era datasets such as Lyman-line and quasar observations (e.g. Sims et al. in prep), as well as deeper JWST data (e.g. $z=15-25$, Pérez-González et al., 2025, published during the final stages of this manuscript). The recent analysis by Zvi Katz et al. (2025), focusing on differentiating Pop II and Pop III stellar population (using HST observations at $z\lesssim 10$, quasar absorption features and CRB measurements), further showcases the exciting prospects that exist in synergies between multi-wavelength observations of the early Universe.
@@ Acknowledgements
JD would like to thank Adam Ormondroyd for his help and useful discussions on PolyChord, and Clàudia Janó Muñoz for comments on the draft.
JD acknowledges support from the Boustany Foundation and Cambridge Commonwealth Trust in the form of an Isaac Newton Studentship. TGJ and EdLA acknowledge the support of the Science and Technology Facilities Council (STFC) through grant numbers ST/V506606/1 and a Rutherford Fellowship, respectively. HTJB would like to acknowledge support from the Kavli Institute for Cosmology Cambridge, and the Kavli Foundation. SP acknowledges the support of the Cambridge Commonwealth, European & International Trust and the Cavendish Laboratory through a Cambridge International & Isaac Newton Studentship. ST acknowledges support by the Royal Society Research Grant G125142. RB acknowledges the support of the Israel Science Foundation (grant No. 2359/20).
The simulations and analysis were performed under DiRAC project number APP15468 using the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1. DiRAC is part of the National e-Infrastructure.
@@
The 30,000 simulations performed using 21cmSPACE, the emulators trained on them, and the PolyChord chains and analysis scripts used in this work can all be made available upon reasonable request to the corresponding author.
@@ Appendix A Impact of fixed UVLF parameters
  • Impact of $t_{\star,\mathrm{II}}$: The effect of the averaging time-scale $t_{\star,\mathrm{II}}$ on the UVLF is to shift the entire distribution by a fixed $\Delta M_{\text{UV}}=2.5\times\Delta\log(t_{\star,\mathrm{II}})$. This is because the UVLF is a direct tracer of the SFR, which is scaled by $t_{\star,\mathrm{II}}$. In Figure 10, we see the effect of varying $t_{\star,\mathrm{II}}=0.2\rightarrow 0.5$ resulting in a shift in magnitude $\Delta M_{\text{UV}}=1\,\mathrm{mag}$. The parameter is completely degenerate with $\tilde{f}_{\star,\mathrm{II}}$, introduced in Equation 7, and can be subsumed into its interpretation (see Footnote 4). Since we typically account for an error magnitude of 20% from semi-numerical modelling in our Bayesian analysis, the effect of $t_{\star,\mathrm{II}}$ (varied by a factor of $2-3$) is within the error budget of our inference.
Appendix B Emulator accuracy
Figure 11: Errors on the 1D emulators for the 21-cm global signal, cosmic X-ray background, cosmic radio background, and 2D emulators for the 21-cm power spectrum and UV luminosity functions, as a function of the input space (e.g., $z$ for the global signal) shown at the 68% and 95% confidence levels calculated using the test data. The third column shows the test data alongside its emulated counterpart, at the 95th percentile of RMSE. The data and shaded regions in yellow show the observational data used in the constraints, for comparison.
Table 6: Summary of works on astrophysical constraints using the code 21cmSPACE. The astrophysical parameters used in each work are listed, along with the observational data used. For details on the parameters, please refer to the cited works — the model changes between works, but the qualitative interpretation of parameters is similar. Previous works use a constant Pop II SFE in the atomic cooling regime ($f_{\star}$ or $f_{\star,\mathrm{II}}$). In this work, we modify and introduce halo mass and redshift-dependent SFE with the efficiency level normalized to $\tilde{f}_{\star,\mathrm{II}}$. We also introduce new parameters $\alpha_{\star}$, $M_{0}$ and $\beta_{\star}$ (see Equation 7). The indicates improved upper limits from HERA Phase 1 (HERA Collaboration et al., 2023), as opposed to the earlier limits (HERA Collaboration et al., 2022b).
@@ Appendix B Emulator accuracy
Pochinda et al. (2024) Multiwavelength synergies $V_{c},f_{\star,\mathrm{II}},f_{\star,\mathrm{III}},t_{\text{delay}},\tau,\alpha_{X},E_{\text{min,X}},f_{X},f_{r}$ HERA P1, SARAS 3, CXB, CRB
@@
where $\theta_{i}$ is the $i$-th parameter set in the test data, $f_{\text{theory/emu}}$ are the theory (simulated) and emulated observables respectively. Instead of performing any averaging, we can sort the errors (over $\theta_{i}$, keeping $x$ bins separate) in ascending order and define 68% and 95% confidence intervals which give us $\epsilon_{1\sigma}(x)$ and $\epsilon_{2\sigma}(x)$ respectively. This gives us an idea of where in the $x$-space the emulator is performing well, and provides a way to accurately account for this error in final analysis. This method has been previously explored in the context of posterior validation for 21-cm emulators in Dorigo Jones et al. (2023) and for CMB emulators in Bolliet et al. (2024). In particular, its utility in our analysis is to check whether the emulator is performing well in the region of the input space we care about the most — i.e., where the observational data lies. For example, in case of the 21-cm signal, we are interested in the SARAS 3 band of $z=15\,$ 2525\,2525. The emulator accuracy outside this range is less important. This can be easily extended to 2D functions like 21-cm power spectrum $\Delta_{21}^{2}(k,z|\theta_{\text{ast}})$ or UVLFs $\Phi(M_{\text{UV}},z|\theta_{\text{ast}})$.
Figure 11 shows the errors on all emulators at the 68% and 95% confidence levels, alongside the observational data. Note that, since the global signal potentially crosses zero, we define the error metric in Equation 39 with $\text{max}(T_{21,\text{true}}(z|\theta_{\text{ast}}))$ instead of $T_{21,\text{true}}(z|\theta_{\text{ast}})$ in the denominator. In principle, this metric can be defined in a variety of ways (using squared error instead of absolute, for example).
@@
$\sigma_{\text{model}}(x)=f(x)\times\sqrt{\epsilon_{\text{theory}}^{2}+\epsilon_{\text{emu},1\sigma}^{2}(x)}\$ (40)
@@ Appendix C Extra tables and figures
Table 6 shows a summary of past works constraining astrophysics using 21cmSPACE  listed for ease of parsing literature and comparing results.
Figure 12 shows the joint constraints on the full parameter space with and without the UVLF dataset, for comparison with the joint analysis in Figure 4. Although the constraints are model dependent, the case without UVLF data is consistent with the inference of Pochinda et al. (2024) who assume a constant SFE in the atomic cooling regime. This is a special case of our more general model, where the SFE depends on both halo-mass and redshift.
@@
Figure 12: Similar to Figure 4, but showing the joint constraints with and without the UVLF dataset. The 2D posterior contours show the 68% (dash-dotted) 95% credible regions (dashed), with filled contours for the case without UVLF data. The 1D marginal posterior PDFs show the weighted mean (solid line) and 68% credible region (shaded with dash-dotted outline). The top right panel shows the prior to posterior volume contraction for each dataset. Note in particular, without the UVLF dataset: $V_{c}$ saturates to upper prior limit, the four SFE parameters are consistent with a constant, fixed value at the lower limit of $0.01\%$ (recovering the model used in Pochinda et al., 2024, and indeed their results), and $f_{X}$ and $f_{r}$ are weakened by their degeneracy with SFE.

### v2

@@ 6 Conclusions
    The star-formation efficiency model, as illustrated in Figure 1, consists of four parameters: $\tilde{f}_{\star,\mathrm{II}}$ scales the magnitude of the SFE, $M_{0}$ sets the halo mass below which the curve flattens, $\alpha_{\star}$ is the power law slope of the $M_{h}$-dependence, and $\beta_{\star}$ controls the redshift-dependence anchored to $z=6$. These parameters are almost entirely constrained by the UVLF data with other data sets providing much weaker constraints. The data favours a non-monotonic evolution of the SFE with redshift: slow evolution at $z\approx 6\,$ 10\,10 with $\beta_{\star}\lesssim 1.4$ (and the posterior saturating to zero, consistent with no evolution), and then a rapid evolution at $z\approx 10-15$ with $\beta_{\star}=3.6^{+0.7}_{-0.7}$ at 68% confidence. The joint fit combining both redshift bins gives a maximum posterior estimate $\beta_{\star}=2.9^{+0.9}_{-0.7}$. Thus, we find consistency with pre-JWST HST-calibrated results at low redshifts (Sun & Furlanetto, 2016; Tacchella et al., 2018), and post-JWST enhanced SFE models at high redshifts (in particular the feedback-free burst model of Dekel et al. 2023; Li et al. 2024, and the density-modulated SFE of Somerville et al. 2025). The slope of the SFE with halo mass $\alpha_{\star}=1.1^{+0.2}_{-0.2}$ is consistent across the two redshift bins for $M_{h}\gtrsim 10^{10}\,\mathrm{M}_{\odot}$, in agreement with the findings of Tacchella et al. (2018). Futhermore, under our composite model, we infer that DM halos of mass $M_{h}=10^{10}\,\mathrm{M}_{\odot}$ have a SFE of 11\, $2\%$ at $z\lesssim 10$, $8\%$ at $z=12$ and $21\%$ at $z=15$ (for the full functional posteriors of the SFE in the low and high-redshift regime, see Figure 7).
@@
    The radio efficiency of early galaxies $f_{r}$, defined relative to those in the local Universe in Equation 19, is constrained by a combination of CRB and HERA data (with a weak contribution from SARAS 3). CRB data disfavours high radio efficiency of galaxies by their direct contribution to the diffuse background seen today. HERA further disfavours strong radio sources since they would boost the amplitude of 21-cm signal fluctuations beyond observed upper limits. The joint analysis results in an upper limit $f_{r}\lesssim 16.9\,(243.7)$ at $68\%$ ($95\%$) confidence levels. We rule out $f_{r}\gtrsim 1800$ at $5\sigma$ confidence level. At $1\sigma$, this is comparable to the constraints obtained by Sikder et al. (2024b) using radio source clustering data from VLA and ATCA (not included here).
@@
    The joint analysis without UVLF data provides weaker constraints on $f_{X}$ and $f_{r}$ due their degeneracy with an unconstrained SFE in the calculation of the diffuse backgrounds (e.g. disfavouring $f_{X}\gtrsim 193$ and $f_{r}\gtrsim 2280$ at $2\sigma$). By including UVLF data and thus informing the model of SFE in observed galaxies, we break this degeneracy and tighten the constraints by a factor of $\sim 10$ (disfavouring $f_{X}\gtrsim 17$ and $f_{r}\gtrsim 244$ at $2\sigma$; see also Zvi Katz et al., 2025, whose result supports this conclusion).
In conclusion, we have shown the critical role of UVLFs in our understanding of the Cosmic Dawn and Epoch of Reionization: they set an upper limit on the minimum mass of star-forming halos, constrain the star-formation efficiency of galaxies across cosmic epochs, and thus also break degeneracies with other astrophysical parameters constrained by different wavelength probes.
[…]
@@ Acknowledgements
JD would like to thank Adam Ormondroyd for his help in understanding and using PolyChord, Will McClymont for discussions on the astrophysics of high-$z$ galaxies, and Clàudia Janó Muñoz for comments on the first draft. JD would also like to thank the anonymous reviewer for their constructive feedback which has improved the manuscript.
JD acknowledges support from the Boustany Foundation and Cambridge Commonwealth Trust in the form of an Isaac Newton Studentship. TGJ and EdLA acknowledge the support of the Science and Technology Facilities Council (STFC) through grant numbers ST/V506606/1 and a Rutherford Fellowship, respectively. HTJB would like to acknowledge support from the Kavli Institute for Cosmology Cambridge, and the Kavli Foundation. SP acknowledges the support of the Cambridge Commonwealth, European & International Trust and the Cavendish Laboratory through a Cambridge International & Isaac Newton Studentship. ST acknowledges support by the Royal Society Research Grant G125142. RB acknowledges the support of the Israel Science Foundation (grant No. 1078/24). The simulations and analysis were performed under DiRAC project number APP15468 using the DiRAC@Durham facility managed by the Institute for Computational Cosmology on behalf of the STFC DiRAC HPC Facility (www.dirac.ac.uk). The equipment was funded by BEIS capital funding via STFC capital grants ST/P002293/1, ST/R002371/1 and ST/S002502/1, Durham University and STFC operations grant ST/R000832/1. DiRAC is part of the National e-Infrastructure.
@@
The 30,000 simulations performed using 21cmSPACE, the trained emulators, the PolyChord chains, and other analysis scripts used in this work can all be made available upon reasonable request to the corresponding author.
@@ Appendix A Impact of fixed UVLF parameters
Figure B1: Errors on the 1D emulators for the 21-cm global signal, cosmic X-ray background, cosmic radio background, and 2D emulators for the 21-cm power spectrum and UV luminosity functions, as a function of the input space (e.g., $z$ for the global signal) shown at the 68% and 95% confidence levels calculated using the test data. The third column shows the test data alongside its emulated counterpart, at the 95th percentile of RMSE. The data and shaded regions in yellow show the observational data used in the constraints, for comparison.
Table C1: Summary of works on astrophysical constraints using the code 21cmSPACE. The astrophysical parameters used in each work are listed, along with the observational data used. For details on the parameters, please refer to the cited works — the model changes between works, but the qualitative interpretation of parameters is similar. Previous works use a constant Pop II SFE in the atomic cooling regime ($f_{\star}$ or $f_{\star,\mathrm{II}}$). In this work, we modify and introduce halo mass and redshift-dependent SFE with the efficiency level normalized to $\tilde{f}_{\star,\mathrm{II}}$. We also introduce new parameters $\alpha_{\star}$, $M_{0}$ and $\beta_{\star}$ (see Equation 7). The indicates improved upper limits from HERA Phase 1 (HERA Collaboration et al., 2023), as opposed to the earlier limits (HERA Collaboration et al., 2022b).
@@
Pochinda et al. (2024) Multi-wavelength synergies $V_{c},f_{\star,\mathrm{II}},f_{\star,\mathrm{III}},t_{\text{delay}},\tau,\alpha_{X},E_{\text{min,X}},f_{X},f_{r}$ HERA P1, SARAS 3, CXB, CRB
@@
  • Impact of $t_{\star,\mathrm{II}}$: The effect of the averaging time-scale $t_{\star,\mathrm{II}}$ on the UVLF is to shift the entire distribution by a fixed $\Delta M_{\text{UV}}=2.5\times\Delta\log(t_{\star,\mathrm{II}})$. This is because the UVLF is a direct tracer of the SFR, which is scaled by $t_{\star,\mathrm{II}}$. In Figure 10, we see the effect of varying $t_{\star,\mathrm{II}}=0.2\rightarrow 0.5$ resulting in a shift in magnitude $\Delta M_{\text{UV}}=1\,\mathrm{mag}$. The parameter is completely degenerate with $\tilde{f}_{\star,\mathrm{II}}$, introduced in Equation 7, and can be subsumed into its interpretation (see Footnote 4). Since we typically account for an error magnitude of 20% from semi-numerical modelling in our Bayesian analysis, the effect of $t_{\star,\mathrm{II}}$ (varied by a factor of 232-3) is within the error budget of our inference.
Appendix B Emulator accuracy
@@
where $\theta_{i}$ is the $i$-th parameter set in the test data, $f_{\text{theory/emu}}$ are the theory (simulated) and emulated observables respectively. Instead of averaging, we can sort the errors (over $\theta_{i}$, keeping $x$ bins separate) in ascending order and define 68% and 95% confidence intervals which yields $\epsilon_{1\sigma}(x)$ and $\epsilon_{2\sigma}(x)$ respectively. This gives us an idea of where in the $x$-space the emulator is performing well, and provides a way to accurately account for this error in final analysis. This method has been previously explored in the context of posterior validation for 21-cm emulators (Dorigo Jones et al., 2023; Breitman et al., 2024) and for CMB emulators (Bolliet et al., 2024). In particular, its utility in our analysis is to check whether the emulator is performing well in the region of the input space we care about the most — i.e., where the observational data lies. For example, in case of the 21-cm signal, we are interested in the SARAS 3 band of z=15z=15\, 25\,25. The emulator accuracy outside this range is less important. This can be easily extended to 2D functions like 21-cm power spectrum $\Delta_{21}^{2}(k,z|\theta_{\text{ast}})$ or UVLFs $\Phi(M_{\text{UV}},z|\theta_{\text{ast}})$.
Figure B1 shows the errors on all emulators at the 68% and 95% confidence levels, alongside the observational data. Note that, since the global signal potentially crosses zero, we define the error metric in Equation 38 with $\text{max}(T_{21,\text{true}}(z|\theta_{\text{ast}}))$ instead of $T_{21,\text{true}}(z|\theta_{\text{ast}})$ in the denominator. In principle, this metric can be defined in a variety of ways (using squared error instead of absolute, for example).
@@
$\sigma_{\text{model}}(x)=f(x)\times\sqrt{\epsilon_{\text{theory}}^{2}+\epsilon_{\text{emu},1\sigma}^{2}(x)}\\$ (39)
@@ Appendix C Extra tables and figures
Table C1 shows a summary of past works constraining astrophysics using 21cmSPACE  listed for ease of parsing literature and comparing results.
Figure 2 shows the joint constraints on the astrophysical parameter space with and without the UVLF dataset, for comparison with the joint analysis in Figure 4. Although the constraints are model dependent, the case without UVLF data is consistent with the inference of Pochinda et al. (2024) who assume a constant SFE in the atomic cooling regime. This is a special case of our more general model, where the SFE depends on both halo-mass and redshift.
Figure 3 shows the joint constraints on the astrophysical parameter space with and without the SARAS 3 dataset, for comparison with the joint analysis in Figure 4. We demonstrate the synergy between SARAS 3 (or more generally, 21-cm global signal limits) and UVLF data in constraining the minimum circular velocity $V_{c}$ for star-formation in DM halos.
@@
Figure 2: Similar to Figure 4, but showing joint constraints with and without the UVLF dataset. The 2D posterior contours show the 68% (dash-dotted) and 95% credible regions (dashed), with filled contours for the case without UVLF data. The 1D marginal posterior PDFs show the weighted mean (solid line) and 68% credible region (shaded with dash-dotted outline). The top right panel shows the prior to posterior volume contraction for each dataset. Note in particular, without the UVLF dataset: $V_{c}$ saturates to upper prior limit, the four SFE parameters are consistent with a constant, fixed value at the lower limit of $0.01\%$ (recovering the model used in Pochinda et al., 2024, and indeed their results), and $f_{X}$ and $f_{r}$ are weakened by their degeneracy with SFE.
Figure 3: Similar to Figure 4, but showing joint constraints with and without the SARAS 3 dataset. The 2D posterior contours show the 68% (dash-dotted) and 95% credible regions (dashed), with filled contours for the case without SARAS 3 data. The 1D marginal posterior PDFs show the weighted mean (solid line) and 68% credible region (shaded with dash-dotted outline). The top right panel shows the prior to posterior volume contraction for each dataset. The SARAS 3 data weakly constrains the minimum circular velocity $V_{c}$ at the lower end of the prior (as opposed to the UVLF data which disfavours the higher end). The constraint is highly sensitive to the SARAS 3 foreground fit as discussed in Section 5.1 (see, e.g. Figure 6 where the constraint is strongest, and the synergy between UVLFs and SARAS 3 is most visible).